# Decoding for Punctured Convolutional and Turbo Codes: A Deep Learning Solution for Protocols Compliance

### v1

@@ Abstract
Neural network-based decoding methods have shown promise in enhancing error correction performance, but traditional approaches struggle with the challenges posed by punctured codes. In particular, these methods fail to address the complexities of variable code rates and the need for protocol compatibility. This paper presents a unified Long Short-Term Memory (LSTM)-based decoding architecture specifically designed to overcome these challenges. The proposed method unifies punctured convolutional and Turbo codes. A puncture embedding mechanism integrates puncturing patterns directly into the network, enabling seamless adaptation to varying code rates, while balanced bit error rate training ensures robustness across different code lengths, rates, and channels, maintaining protocol flexibility. Extensive simulations in Additive White Gaussian Noise and Rayleigh fading channels demonstrate that the proposed approach outperforms conventional decoding techniques, providing significant improvements in decoding accuracy and robustness. These results underscore the potential of LSTM-based decoding as a promising solution for next-generation artificial intelligence powered communication systems.
@@ 1.1 Prior Works
In wireless communication, AI has shown promise in physical layer signal processing tasks such as channel estimation, signal detection, and channel decoding [4, 5, 6, 7, 8, 9]. Traditionally, signal processing algorithms were implemented on Central Processing Units (CPUs), Digital Signal Processors (DSPs), or Application-Specific Integrated Circuits (ASICs), which rely on serial processing. However, the highly parallel nature of AI algorithms requires more efficient architectures, driving the shift from CPU-based signal processing algorithms to Graphics Processing Units (GPUs), which are optimized for parallel computation and can efficiently handle large-scale, simultaneous operations.
@@
Puncturing, which discards part of the encoded data to form different code rates and improve spectral efficiency, is essential in real-world wireless communication systems. Control channels typically employ lower code rates for high reliability, while data channels use more flexible and higher code rates to accommodate diverse transmission conditions and maximize throughput. Linear block codes (e.g., Low-Density Parity-Check (LDPC) and Polar codes) and sequential codes (e.g., convolutional and Turbo codes) are widely used in commercial communication protocols [14, 15, 16, 17]. For example, Wi-Fi protocols [18] support four code rates for convolutional and LDPC codes, while cellular protocols [19, 20] utilize Polar codes for control channels with dozens of code rates, and Turbo and LDPC codes for data channels with over a hundred code rates.
@@
Other studies have focused on applying neural networks to decode sequential codes, such as convolutional and Turbo codes. For example, [23] explored the use of Recurrent Neural Networks (RNNs) for decoding convolutional codes, achieving performance comparable to the Viterbi algorithm [24] in both AWGN and non-AWGN channels with $t$-distributed noise. The DeepTurbo approach [25], inspired by the iterative Bahl-Cocke-Jelinek-Raviv (BCJR) algorithm [26], targets Turbo codes but faces performance degradation when generalized to longer code lengths, requiring retraining. A notable study [27] employed model-agnostic meta-learning to enhance the generalization of neural network-based decoders in unseen channel conditions. Additionally, Turbo Autoencoders [28] introduced an end-to-end learning framework that jointly optimizes both the encoder and decoder, outperforming traditional BCJR decoders.
@@ 1.2 Contributions
To address this gap of puncturing issues in sequential codes, we propose a unified, protocol-compatible Long Short-Term Memory (LSTM)-based neural decoder for punctured convolutional and Turbo codes, inspired by the inherent similarity between convolutional coding, which transmits information through state transitions in shift registers, and the information propagation across time steps in LSTM networks. The proposed architecture encodes puncturing patterns directly into the neural network’s latent space and introduces a Balanced Bit Error Rate Training (BBT) method for efficient fine-tuning across different code rates, ensuring seamless compatibility with protocol flexibility  [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html. The main contributions of this work are summarized as follows.
@@
    Protocol-Compatible Neural Decoding: We propose a unified protocol-compatible neural decoding framework specifically designed for punctured convolutional and Turbo codes. This framework ensures compliance with Wi-Fi (IEEE 802.11) and cellular (3GPP TS 36.212) standards by supporting varying code lengths and rates, even under practical puncturing patterns. The protocol compatibility is achieved through puncture embedding and balanced bit error rate training, which enables the decoder to handle different code rates effectively. To the best of our knowledge, this is the first neural decoding approach capable of generalizing seamlessly across diverse protocol-compliant configurations while maintaining competitive performance.
@@
    Puncture Embedding for Adaptability: We introduce a puncture embedding module that encodes puncturing patterns directly into the neural network’s latent space, enabling seamless adaptation to various code rates. By incorporating the puncture embedding as a gating mechanism, the flow of Log-Likelihood Ratio information is controlled. This design allows the decoder to support diverse code lengths and rates with a single set of network parameters, ensuring exceptional generalization and compatibility with dynamically changing protocol requirements, making it highly suitable for modern wireless communication systems where adaptability is critical.
@@
    Superior Performance in Practical Channels: The proposed neural decoder demonstrates state-of-the-art performance in both Additive White Gaussian Noise and Rayleigh fading channels. Extensive simulations show that the model not only outperforms traditional decoding algorithms under practical Least Squares (LS) channel estimation but also exceeds the performance of conventional decoders with Perfect Channel State Information (PCSI).
This paper is structured as follows. Section II offers foundational knowledge, covering the basics of convolutional codes, Turbo codes, and Long Short-Term Memory networks. Section III describes the proposed Convolutional Neural Engine in detail. Section IV discusses the outcomes of training and inference experiments. Section V evaluates the computational complexity and decoding latency of the proposed architecture. Finally, Section VI provides a conclusion.
Notations: Bold lowercase and uppercase letters are used to represent vectors and matrices, respectively. The symbol $\mathbb{R}$ represents the set of real numbers, while $\mathbb{C}$ denotes the set of complex numbers. The transpose operation is denoted by $[\cdot]^{T}$, and $[\cdot]^{H}$ represents the Hermitian (conjugate transpose) operation. The notation $\mathcal{N}(\mu,{\sigma}^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance ${\sigma}^{2}$.
@@ 2.1 Fundamentals of Convolutional and Turbo Codes
The convolutional code is characterized by its encoder structure, where the input bit stream passes through a series of shift registers, generating encoded bits as output. Each output bit is determined by the current input bit and several previous input bits, as defined by the constraint length of the code. The constraint length represents the number of bits stored in the encoder’s memory that influence the current output. In IEEE 802.11, the convolutional encoder operates with a mother code rate of 1/2 and a constraint length of 7. This setup produces two output bits for each input bit, with the constraint length indicating that the output depends on the current bit and the six preceding bits. Figure 1 provides a schematic illustration of the convolutional encoder used in IEEE 802.11, showcasing its encoding process and structure. The bit denoted as "A" shall be output from the encoder before the bit denoted as "B". The specific generator polynomials governing this encoding process, expressed in octal form, are provided in Eq. 1. Higher code rates, such as 2/3, 3/4, and 5/6, are achieved using a technique known as puncturing. Puncturing selectively removes specific encoded bits to increase the effective code rate without altering the encoder’s fundamental structure. Figure 2 illustrates how puncturing is applied in IEEE 802.11 to achieve these higher code rates.
@@
In 3GPP TS 36.212, Turbo codes are used in the downlink and uplink channels for the physical layer, where they provide strong error correction capabilities. The standard adopts Parallel Concatenated Convolutional Codes (PCCC) as the foundation of Turbo coding. Figure 3 illustrates the encoding process of Turbo codes, which consists of two parallel convolutional encoders, each with a constraint length of 5. The first encoder processes the input bitstream directly, while the second operates on an interleaved version of the input, with the interleaver introducing randomness to the input sequence. The specific generator polynomials for these encoders are detailed in Equation 2.
@@ 2.2 Long Short-Term Memory Neural Network
Long Short-Term Memory networks [29], a variant of Recurrent Neural Networks [30], are specifically designed to model sequential data and capture long-term dependencies effectively. As shown in Figure 5, RNNs employ a cyclic structure that allows information to persist across time steps $t$, enabling sequential data processing. The update of the hidden state $h_{t}$ in an RNN is governed by:
$\boldsymbol{h}_{t}=\tanh\left(\boldsymbol{W}_{h}\boldsymbol{h}_{t-1}+\boldsymbol{W}_{x}\boldsymbol{x}_{t}+\boldsymbol{b}_{h}\right)$ (3)
where $\boldsymbol{h}_{t}$ is the hidden state at time step $t$, $\boldsymbol{h}_{t-1}$ is the hidden state from the previous time step, $\boldsymbol{x}_{t}$ is the input at time step $t$, and $\boldsymbol{W}_{h}$, $\boldsymbol{W}_{x}$, $\boldsymbol{b}_{h}$ are learnable parameters. This cyclic nature of RNNs makes them well-suited for capturing temporal dependencies in sequential data. However, traditional RNNs encounter challenges during training, particularly vanishing and exploding gradient issues, which hinder their ability to capture long-range dependencies.
@@
$\begin{array}[]{l}\boldsymbol{f}_{t}=\sigma\left(\boldsymbol{W}_{f}\cdot\begin{bmatrix}\boldsymbol{h}_{t-1},\boldsymbol{x}_{t}\end{bmatrix}+\boldsymbol{b}_{f}\right)\\ \boldsymbol{i}_{t}=\sigma\left(\boldsymbol{W}_{i}\cdot\begin{bmatrix}\boldsymbol{h}_{t-1},\boldsymbol{x}_{t}\end{bmatrix}+\boldsymbol{b}_{i}\right)\\ \boldsymbol{o}_{t}=\sigma\left(\boldsymbol{W}_{o}\cdot\begin{bmatrix}\boldsymbol{h}_{t-1},\boldsymbol{x}_{t}\end{bmatrix}+\boldsymbol{b}_{o}\right)\\ \tilde{\boldsymbol{c}}_{t}=\tanh\left(\boldsymbol{W}_{c}\cdot\begin{bmatrix}\boldsymbol{h}_{t-1},\boldsymbol{x}_{t}\end{bmatrix}+\boldsymbol{b}_{c}\right)\\ \boldsymbol{c}_{t}=\boldsymbol{f}_{t}\odot\boldsymbol{c}_{t-1}+\boldsymbol{i}_{t}\odot\tilde{\boldsymbol{c}}_{t}\\ \boldsymbol{h}_{t}=\boldsymbol{o}_{t}\odot\tanh\left(\boldsymbol{c}_{t}\right)\end{array}$ (4)
where $\boldsymbol{f}_{t}$, $\boldsymbol{i}_{t}$, and $\boldsymbol{o}_{t}$ represent the forget, input, and output gate activations at time step $t$; $\boldsymbol{x}_{t}$ is the input; $\tilde{\boldsymbol{c}}_{t}$ is the candidate memory cell; $\boldsymbol{c}_{t}$ is the cell state; and $\boldsymbol{h}_{t}$ is the hidden state.
Here, $\sigma$ denotes the sigmoid activation function, which plays a critical gating role by producing values in the range of $[0,1]$. This enables $\sigma$ to act as a "soft switch" that determines the degree to which information should be passed through the network. For instance, in the forget gate $\boldsymbol{f}_{t}$, $\sigma$ dynamically adjusts the proportion of the previous cell state $\boldsymbol{c}_{t-1}$ that should be retained. Similarly, in the input gate $\boldsymbol{i}_{t}$ and output gate $\boldsymbol{o}_{t}$, $\sigma$ controls how much new information is incorporated into the cell state and how much of the cell state influences the hidden state $\boldsymbol{h}_{t}$, respectively. By gating these information flows, $\sigma$ ensures that the network can focus on relevant inputs and maintain stable gradients during training, allowing LSTMs to effectively capture both short- and long-term dependencies.
@@ 2.3 System Model
The proposed system is evaluated under two distinct communication channel models: the AWGN channel and the Rayleigh fading channel. The AWGN channel serves as the training and inference environment, providing an idealized setting for model training. In contrast, the Rayleigh fading channel is used solely for inference, allowing us to assess how well the model, trained in the AWGN channel, generalizes to more realistic fading conditions.
In the AWGN channel, the communication link follows the sequence illustrated in Figure 7, where $K$ denotes the length of the original message, $N$ represents the length of the encoded message, and $E$ refers to the length of the message after rate-matching. The code rate $R$ of the encoded message is defined as the ratio of the message length to the rate-matched length, i.e., $R=\frac{K}{E}$.
@@
The transmitted signal $\boldsymbol{x}$ is modulated using Binary Phase Shift Keying (BPSK), and the received signal $\boldsymbol{y}$ is corrupted by additive Gaussian noise $\boldsymbol{n}$. The relationship between the transmitted and received signals is expressed as:
$\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{n}$ (5)
where $\boldsymbol{n}\sim\mathcal{N}(0,\sigma^{2})$ represents the AWGN noise. After demodulation and de-rate matching, the signal is decoded to obtain the transmitted message.
@@
$\boldsymbol{y}=\boldsymbol{H}\boldsymbol{x}+\boldsymbol{n}$ (6)
where $\boldsymbol{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is the channel matrix representing the Rayleigh fading coefficients, with $N_{R}$ denoting the number of receive antennas and $N_{T}$ the number of transmit antennas. $\boldsymbol{x}\in\mathbb{C}^{N_{T}\times 1}$ is the transmitted signal vector in the frequency domain after modulation (e.g., OFDM subcarriers). $\boldsymbol{n}\in\mathbb{C}^{N_{R}\times 1}$ is the noise vector at the receiver, where each element is i.i.d. Gaussian noise with zero mean and variance $\sigma^{2}$.
The channel matrix $\boldsymbol{H}$ is constructed by generating multipath coefficients in the time domain, based on an $L$-tap delay profile determined by environmental characteristics and delay spread, and then transformed into the frequency domain via the Fast Fourier Transform (FFT) to be applied to $\boldsymbol{x}$. Each tap of $\boldsymbol{H}$ is modeled as an independent complex Gaussian random variable, and the total power of the $L$ taps is normalized to 1, i.e., $\sum_{l=0}^{L-1}\mathbb{E}[|h_{l}|^{2}]=1$, where $h_{l}$ represents the coefficient of the $l$-th tap.
Channel estimation is performed to estimate $\boldsymbol{H}$, and Multiple Input Multiple Output (MIMO) detection is used to recover the transmitted signal $\boldsymbol{x}$ after it has been affected by multipath fading. The subsequent steps, including demodulation, de-rate matching, and decoding, are similar to those in the AWGN case, with the goal of recovering the transmitted message.

### v3

@@ Abstract
Neural network-based decoding methods show promise in enhancing error correction performance but face challenges with punctured codes. In particular, existing methods struggle to adapt to variable code rates or meet protocol compatibility requirements. This paper proposes a unified long short-term memory (LSTM)-based neural decoder for punctured convolutional and Turbo codes to address these challenges. The key component of the proposed LSTM-based neural decoder is puncturing-aware embedding, which integrates puncturing patterns directly into the neural network to enable seamless adaptation to different code rates. Moreover, a balanced bit error rate training strategy is designed to ensure the decoder’s robustness across various code lengths, rates, and channels. In this way, the protocol compatibility requirement can be realized. Extensive simulations in both additive white Gaussian noise (AWGN) and Rayleigh fading channels demonstrate that the proposed neural decoder outperforms conventional decoding techniques, offering significant improvements in decoding accuracy and robustness.
@@ 1.1 Prior Works
Building on AI’s potential in telecommunications, its application in physical layer signal processing for wireless communications has garnered significant attention, particularly in tasks such as channel estimation, signal detection, and channel decoding [4, 5, 6, 7, 8, 9], where it demonstrates notable advantages. In contrast, traditional signal processing algorithms have typically been implemented on central processing units (CPUs), digital signal processors (DSPs), or application-specific integrated circuits (ASICs), relying on serial processing. However, due to the highly parallel nature of AI algorithms, traditional serial processing architectures struggle to meet their demands, thus driving the shift toward graphics processing units (GPUs), which are optimized for parallel computation and capable of efficiently handling large-scale, simultaneous operations.
@@
Puncturing, which discards part of the encoded data to form different code rates and improve spectral efficiency, is essential in real-world wireless communication systems. Control channels typically employ lower code rates for high reliability, while data channels use more flexible and higher code rates to accommodate diverse transmission conditions and maximize throughput. Linear block codes (e.g., low-density parity-check (LDPC) and Polar codes) and sequential codes (e.g., convolutional and Turbo codes) are widely used in commercial communication protocols [14, 15, 16, 17]. For example, Wi-Fi protocols [18] support four code rates for convolutional and LDPC codes, while cellular protocols [19, 20] utilize Polar codes for control channels with dozens of code rates, and Turbo and LDPC codes for data channels with over a hundred code rates. However, the puncturing arrangements in these standards may not always be optimal, as they are often designed for implementation convenience rather than performance maximization. Extrinsic information transfer (EXIT) charts provide a powerful semi-analytical tool for designing and analyzing iteratively decoded systems, including the optimization of puncturing schemes for Turbo codes [21, 22, 23, 24].
@@
Other studies have focused on applying neural networks to decode sequential codes, such as convolutional and Turbo codes. For example, [27] explored the use of recurrent neural networks (RNNs) for decoding convolutional codes, achieving performance comparable to the Viterbi algorithm [28] in both AWGN and non-AWGN channels with $t$-distributed noise. The DeepTurbo approach [29], inspired by the iterative Bahl-Cocke-Jelinek-Raviv (BCJR) algorithm [30], targets Turbo codes but faces performance degradation when generalized to longer code lengths, requiring retraining. A notable study [31] employed model-agnostic meta-learning to enhance the generalization of neural network-based decoders in unseen channel conditions. Additionally, Turbo autoencoders [32] introduced an end-to-end learning framework that jointly optimizes both the encoder and decoder, outperforming traditional BCJR decoders under canonical channels.
@@ 1.2 Contributions
To address this gap of puncturing issues in sequential codes, we propose a unified, protocol-compatible long short-term memory (LSTM)-based neural decoder for punctured convolutional and Turbo codes. This approach is inspired by the similarity between the LSTM’s memory mechanism and the shift register’s role in convolutional codes, as both retain a history of prior inputs to inform current outputs. The proposed approach encodes puncturing patterns directly into the neural network’s latent space and introduces a balanced bit error rate training (BBT) method for efficient fine-tuning across different code rates, ensuring seamless compatibility with protocol flexibility [^1] Simulation codes will be provided to reproduce the results in this paper: http://oa.ee.tsinghua.edu.cn/dailinglong/publications/publications.html.. The main contributions of this work are summarized as follows.
@@
    Protocol-Compatible Neural Decoding: We propose a unified protocol-compatible neural decoding approach specifically designed for punctured convolutional and Turbo codes. This approach ensures compliance with Wi-Fi (IEEE 802.11) and cellular (3GPP TS 36.212) standards by supporting varying code lengths and rates, even under practical puncturing patterns. The protocol compatibility is achieved through puncturing-aware embedding and balanced bit error rate training, which enables the decoder to handle different code rates effectively. To the best of our knowledge, this is the first neural decoding approach capable of generalizing seamlessly across diverse protocol-compliant configurations while maintaining competitive performance.
@@
    Puncturing-Aware Embedding for Adaptability: We introduce a puncturing-aware embedding module that encodes puncturing patterns directly into the neural network’s latent space, enabling seamless adaptation to various code rates. By incorporating the puncturing-aware embedding as a gating mechanism, the flow of log-likelihood ratio information is controlled. This design allows the decoder to support diverse code lengths and rates with a single set of network parameters, ensuring exceptional generalization and compatibility with dynamically changing protocol requirements, making it highly suitable for modern wireless communication systems where adaptability is critical.
@@
    Superior Performance in Practical Channels: The proposed neural decoder demonstrates state-of-the-art performance in both additive white Gaussian noise (AWGN) and Rayleigh fading channels. Extensive simulations show that the model not only outperforms traditional decoding algorithms under practical least squares (LS) channel estimation but also exceeds the performance of conventional decoders with perfect channel state information (PCSI). Moreover, under matched code length and rate conditions during training and inference, the proposed neural decoder yields a 0.2 dB performance gain over DeepTurbo [29] at a BER of $10^{-4}$. In mismatched conditions, it exhibits substantial performance improvements over DeepTurbo.
This paper is structured as follows. Section II offers foundational knowledge, covering the basics of convolutional codes, Turbo codes, and long short-term memory networks. Section III describes the proposed convolutional neural engine in detail. Section IV discusses the outcomes of training and inference experiments. Section V evaluates the computational complexity and decoding latency of the proposed approach. Finally, Section VI provides a conclusion.
Notations: Bold lowercase and uppercase letters are used to represent vectors and matrices, respectively. The symbol $\mathbb{R}$ represents the set of real numbers, while $\mathbb{C}$ denotes the set of complex numbers. The transpose operation is denoted by $\cdot]^{T}$, and $\cdot]^{H}$ represents the Hermitian (conjugate transpose) operation. The notation $\mathcal{N}(\mu,{\sigma}^{2})$ denotes a Gaussian distribution with mean $\mu$ and variance ${\sigma}^{2}$.
@@ 2.1 Fundamentals of Convolutional and Turbo Codes
The convolutional code is characterized by its encoder structure, where the input bit stream passes through a series of shift registers, generating encoded bits as output. Each output bit is determined by the current input bit and several previous input bits, as defined by the constraint length of the code. The constraint length represents the number of bits stored in the encoder’s memory that influence the current output. In IEEE 802.11, the convolutional encoder operates with a mother code rate of 1/2 and a constraint length of 7. This setup produces two output bits for each input bit, with the constraint length indicating that the output depends on the current bit and the six preceding bits. Figure 1 provides a schematic illustration of the convolutional encoder used in IEEE 802.11, showcasing its encoding process and structure. The bit denoted as “zz” shall be output from the encoder before the bit denoted as $z^{\prime}$”. The specific generator polynomials governing this encoding process, expressed in octal form, are provided in Eq. 1. Higher code rates, such as 2/3, 3/4, and 5/6, are achieved using a technique known as puncturing. Puncturing selectively removes specific encoded bits to increase the effective code rate without altering the encoder’s fundamental structure. Figure 2 illustrates how puncturing is applied in IEEE 802.11 to achieve these higher code rates.
@@
In 3GPP TS 36.212, Turbo codes are used in the downlink and uplink channels for the physical layer, where they provide strong error correction capabilities. The standard adopts parallel concatenated convolutional codes (PCCC) as the foundation of Turbo coding. Figure 3 illustrates the encoding process of Turbo codes, which consists of two parallel convolutional encoders, each with a constraint length of 4. The first encoder processes the input bitstream directly, while the second operates on an interleaved version of the input, with the interleaver introducing randomness to the input sequence. The specific generator polynomials for these encoders are detailed in Equation 2.
@@ 2.2 Long Short-Term Memory Neural Network
Long short-term memory networks [33], a variant of recurrent neural networks [34], are specifically designed to model sequential data and capture long-term dependencies effectively. As shown in Figure 5, RNNs employ a cyclic structure that allows information to persist across time steps $t$, enabling sequential data processing. The update of the hidden state $\bm{h}_{t}$ in an RNN is governed by:
$\bm{h}_{t}=\tanh\left(\bm{W}_{h}\bm{h}_{t-1}+\bm{W}_{x}\bm{x}_{t}+\bm{b}_{h}\right)$ (3)
where $\bm{h}_{t}$ is the hidden state at time step $t$, $\bm{h}_{t-1}$ is the hidden state from the previous time step, $\bm{x}_{t}$ is the input at time step $t$, and $\bm{W}_{h}$, $\bm{W}_{x}$, $\bm{b}_{h}$ are learnable parameters. This cyclic nature of RNNs makes them well-suited for capturing temporal dependencies in sequential data. However, traditional RNNs encounter challenges during training, particularly vanishing and exploding gradient issues, which hinder their ability to capture long-range dependencies.
@@
$\begin{array}[]{l}\bm{f}_{t}=\phi\left(\bm{W}_{f}\cdot\begin{bmatrix}\bm{h}_{t-1},\bm{x}_{t}\end{bmatrix}+\bm{b}_{f}\right)\\ \bm{i}_{t}=\phi\left(\bm{W}_{i}\cdot\begin{bmatrix}\bm{h}_{t-1},\bm{x}_{t}\end{bmatrix}+\bm{b}_{i}\right)\\ \bm{o}_{t}=\phi\left(\bm{W}_{o}\cdot\begin{bmatrix}\bm{h}_{t-1},\bm{x}_{t}\end{bmatrix}+\bm{b}_{o}\right)\\ \tilde{\bm{c}}_{t}=\tanh\left(\bm{W}_{c}\cdot\begin{bmatrix}\bm{h}_{t-1},\bm{x}_{t}\end{bmatrix}+\bm{b}_{c}\right)\\ \bm{c}_{t}=\bm{f}_{t}\odot\bm{c}_{t-1}+\bm{i}_{t}\odot\tilde{\bm{c}}_{t}\\ \bm{h}_{t}=\bm{o}_{t}\odot\tanh\left(\bm{c}_{t}\right)\end{array}$ (4)
where $\bm{f}_{t}$, $\bm{i}_{t}$, and $\bm{o}_{t}$ represent the forget, input, and output gate activations at time step $t$; $\bm{x}_{t}$ is the input; $\tilde{\bm{c}}_{t}$ is the candidate memory cell; $\bm{c}_{t}$ is the cell state; and $\bm{h}_{t}$ is the hidden state.
Here, $\phi$ denotes the sigmoid activation function, which plays a critical gating role by producing values in the range of $[0,1]$. This enables $\phi$ to act as a “soft switch” that determines the degree to which information should be passed through the network. For instance, in the forget gate $\bm{f}_{t}$, $\phi$ dynamically adjusts the proportion of the previous cell state $\bm{c}_{t-1}$ that should be retained. Similarly, in the input gate $\bm{i}_{t}$ and output gate $\bm{o}_{t}$, $\phi$ controls how much new information is incorporated into the cell state and how much of the cell state influences the hidden state $\bm{h}_{t}$, respectively. By gating these information flows, $\phi$ ensures that the network can focus on relevant inputs and maintain stable gradients during training, allowing LSTMs to effectively capture both short- and long-term dependencies.
@@ 2.3 System Model
The proposed approach is evaluated under two distinct communication channel models: the AWGN channel and the Rayleigh fading channel. The AWGN channel serves as the training and inference environment, providing an idealized setting for model training. In contrast, the Rayleigh fading channel is used solely for inference, allowing us to assess how well the model, trained in the AWGN channel, generalizes to more realistic fading conditions.
In the AWGN channel, the communication link follows the sequence illustrated in Figure 7, where $K$ denotes the length of the original message, $N$ represents the length of the encoded message, and $E$ refers to the length of the message after rate-matching. The code rate $R$ is defined as the ratio of the original message length to the rate-matched length, i.e., $R=\frac{K}{E}$.
@@
The rate-matched message $\bm{m}\in\{0,1\}^{E}$ is modulated using binary phase shift keying (BPSK) to produce the transmitted signal $\bm{x}\in\{\pm 1\}^{E}$. The received signal $\bm{y}\in\mathbb{R}^{E}$ is corrupted by additive Gaussian noise $\bm{n}\in\mathbb{R}^{E}$, and the relationship between the transmitted and received signals is expressed as:
$\bm{y}=\bm{x}+\bm{n}$ (5)
where $\bm{n}\sim\mathcal{N}(0,\sigma^{2})$ represents the AWGN noise. After demodulation and de-rate matching, the signal is decoded to obtain the original message $\bm{u}\in\{0,1\}^{K}$.
@@
$\bm{y}=\bm{H}\bm{x}+\bm{n}$ (6)
where $\bm{H}\in\mathbb{C}^{N_{R}\times N_{T}}$ is the channel matrix representing the Rayleigh fading coefficients, with $N_{R}$ denoting the number of receive antennas and $N_{T}$ the number of transmit antennas. $\bm{x}\in\mathbb{C}^{N_{T}\times 1}$ is the transmitted signal vector in the frequency domain after modulation (e.g., OFDM subcarriers). $\bm{n}\in\mathbb{C}^{N_{R}\times 1}$ is the noise vector at the receiver, where each element is i.i.d. Gaussian noise with zero mean and variance $\sigma^{2}$.
The channel matrix $\bm{H}$ is constructed by generating multipath coefficients in the time domain, based on an $L$-tap delay profile determined by environmental characteristics and delay spread, and then transformed into the frequency domain via the fast fourier transform (FFT) to be applied to $\bm{x}$. Each tap of $\bm{H}$ is modeled as an independent complex Gaussian random variable, and the total power of the $L$ taps is normalized to 1, i.e., $\sum_{l=0}^{L-1}\mathbb{E}[|h_{l}|^{2}]=1$, where $h_{l}$ represents the coefficient of the $l$-th tap.
Channel estimation is performed to estimate $\bm{H}$, and multiple input multiple output (MIMO) detection is used to obtain the estimated transmitted signal $\hat{\bm{x}}$, after $\bm{x}$ has been affected by multipath fading. The subsequent steps, including demodulation, de-rate matching, and decoding, are similar to those in the AWGN case, with the goal of recovering the original message.